# New Obstacles to Multiple Recurrence

### v2

@@ 1 Introduction
It was Katznelson [17] who proved the equivalence of these two definitions of topological recurrence, giving the eponymous conjecture its combinatorial form. One direction is relatively simple. If a set $S$ is not a set of dynamical topological recurrence, then one can use the dynamical system to come up with a coloring of the Cayley graph. In particular there will be some $\varepsilon$ so that $d(T^{n}x,x)>\varepsilon$ for all $n\in S$. Now cover the compact space by $\varepsilon$-balls for a suitably small $\varepsilon$, and color $m\in\mathbb{N}$ by which $\varepsilon$-ball $T^{m}(x)$ ends up in. The other direction, using a coloring to construct a dynamical system, follows from an argument of Katznelson that makes use of symbolic dynamics, and bears some resemblance to the Furstenberg correspondence principle.
@@
Conjecture 1.6 (Katznelson’s question).
@@
No discussion of the history of Katznelson’s question would be complete without also referencing the “Related Problem”. Bergelson asked whether there is a set which is a set of topological recurrence but not a set of measurable recurrence, or equivalently whether Cayley graph on the integers with a positive upper density independent set, but infinite chromatic number, and Kříž [18] resolved this question in the negative. His construction is based on the boolean Cayley graph over $(\mathbb{Z}/2)^{n}$ generated by elements of Hamming weight $n-10\sqrt{n}$. One can check that elements of Hamming weight $\leq n/2-10\sqrt{n}$ form a positive density independent set in this graph. Furthermore, looking at the elements of Hamming weight $n/2-6\sqrt{n}$, this graph has a copy of the Kneser graph $KG(n,n/2-6\sqrt{n})$ and thus has chromatic number at least $\sqrt{n}$, and goes to infinity with $n$. Kříž projects this construction to the integers. See [12] for an exposition of his example by Griesmer and [11] for an extension.
In this paper, we will be concerned with the natural higher-order version of Katznelson’s question. This was first explicitly asked by Huang, Shao, and Ye in 2011 ([15], [16]), and asked as Question 7 in [7]. Before stating it, we need to define topological multiple recurrence.
@@
The set of $n$ so that these generalized polynomials have small $\mathbb{R}/\mathbb{Z}$ is called a “Nil-Bohr” set, or a $Nil_{d}$-Bohr” set when we want to emphasize the parameter of $d$. An ordinary Bohr set is just a $Nil_{1}$ Bohr set.
@@ 2.1 Building the $\ell^{1}$ example
Let $\delta_{1}$ and $\delta_{2}$ be small positive numbers so that $\delta_{1}/\delta_{2}$ is an integer and $\delta_{2}$ is a suitably small function of $\delta_{1}$, so for instance $\delta_{2}<\delta_{1}^{3}$. For instance one can pick $\delta_{1}=10^{-2}$ and $\delta_{2}=10^{-10}$. However, for ease of notation we will write $\delta_{1}$ and $\delta_{2}$.
@@
Let $m\in\mathbb{N}\cup\infty$. We say that $B\subset\ell^{1}(\mathbb{N}_{\leq m};\mathbb{R}/\mathbb{Z})$ is a Bohr set if there exists $k$ and $\varepsilon$ and a linear function $f:\ell^{1}(\mathbb{N}_{\leq m};\mathbb{R}/\mathbb{Z})\to(\mathbb{R}/\mathbb{Z})^{k}$ so that for $B$ is the set of $b$ so that $|f(b)|<\varepsilon$.
In other words, we must check that for all linear function $f:\ell^{1}(\mathbb{N};\mathbb{R}/\mathbb{Z})\to(\mathbb{R}/\mathbb{Z})^{k}$, there are points of $S$ that are mapped arbitrarily close to the origin. What we actually need is a finitary version of this.
@@
In fact we only used here that the function is additive, since $\delta_{1}/\delta_{2}\in\mathbb{N}$. The astute reader may notice that the linear functions $f:\ell^{1}(\mathbb{N};\mathbb{R}/\mathbb{Z})\to\mathbb{R}/\mathbb{Z}$ comprise the dual group $\ell^{\infty}(\mathbb{N};\mathbb{R}\/\mathbb{Z})$, are given by a sequence of integers $(b_{1},b_{2},\cdots)$ with some $C$ so that $|b_{i}|\leq C$ for all $i$, and send $(a_{1},a_{2},\cdots)$ to $\sum_{i\in\mathbb{N}}a_{i}b_{i}$. In particular, a function to $(\mathbb{R}/\mathbb{Z})^{k}$ will consist of $k$ such functions, one for each coordinate.
@@ 2.2 The coloring
Definition 2.5 (Our Coloring).
Partition $\mathbb{C}/\mathbb{Z}[i]\cong(\mathbb{R}/\mathbb{Z})^{2}$ into $O(1/\delta^{2}_{2})$ sets of diameter at most $\delta_{2}$. Color $(a_{1},\cdots,a_{m})\in\ell^{1}(\mathbb{N}_{\leq m};\mathbb{R}/\mathbb{Z})$ by the set that contains $\sum_{i=1}^{n}(e(a_{i})-1)$ taken modulo $\mathbb{Z}[i]$.
@@ 3 Concluding Remarks and Open Questions
We generally believe that if Katznelson’s question has a positive answer, the proof somehow relies on topology. Indeed Glasscock, Koutsogiannis, and Richter [7] prove some special cases by using the intermediate value theorem, and [18] relies on the topological lower bound for the chromatic number of the Kneser graph.
We mention one important obstacle to extending something like our construction to settle the full Katznelson’s question. Let $G$ be an Abelian group. Then an element of $G^{m}$ in the generating set with sum of coordinates 0 can be written in the form $(a_{1}-a_{2},a_{2}-a_{3},\cdots,a_{m-1}-a_{m},a_{m}-a_{1})$, so the cyclic shifts $(a_{1},a_{2},a_{3},\cdots,a_{n})$ and $(a_{2},a_{3},\cdots,a_{n},a_{1})$. Hence, we cannot simply take a function of each coordinate and add it up the way we considered (e(x)1)(e(x)-1) here. Unfortunately, this kind of “cyclic shift” obstacle destroys many potential counterexamples to Katznelson’s question, and any kind of high-dimensional projection must break this sort of symmetry.
@@
For an Abelian Cayley graph, we can determine in polynomial time whether there exists a coloring by Bohr sets using <C<C colors by simply checking all such colorings, and in [17] it is proved that Katznelson’s Question is equivalent to the statement that the number of colors necessary in such a coloring is bounded a function of the chromatic number. Hence, if Katznelson’s Question has a positive answer, there is a polynomial algorithm for this “Approximate Graph Coloring” question for Abelian Cayley graphs. For general graphs, such problems are known to be NP-hard conditional on the Unique Games Conjecture, which gives us reason to believe this conjecture. See [5] and the references there for more discussion of Approximate Graph Coloring.
@@
For what sets $S$ does a finite coloring of $\mathbb{N}$ contain a monochromatic $\{x,x+d,x+d^{2}\}$ with $d\in S$?

### v3

@@ 1 Introduction
It was Katznelson [17] who proved the equivalence of these two definitions of topological recurrence, giving the eponymous question its combinatorial form. One direction is relatively simple. If a set $S$ is not a set of dynamical topological recurrence, then one can use the dynamical system to come up with a coloring of the Cayley graph. In particular there will be some $\varepsilon$ so that $d(T^{n}x,x)>\varepsilon$ for all $n\in S$. Now cover the compact space by $\varepsilon$-balls for a suitably small $\varepsilon$, and color $m\in\mathbb{N}$ by which $\varepsilon$-ball $T^{m}(x)$ ends up in. The other direction, using a coloring to construct a dynamical system, follows from an argument of Katznelson that makes use of symbolic dynamics, and bears some resemblance to the Furstenberg correspondence principle.
@@
Question 1.6 (Katznelson’s question).
@@
No discussion of the history of Katznelson’s question would be complete without also referencing the “Related Problem”. Bergelson asked whether there is a set which is a set of topological recurrence but not a set of measurable recurrence, or equivalently whether Cayley graph on the integers with a positive upper density independent set, but infinite chromatic number, and Kříž [18] resolved this question in the negative. His construction is based on the boolean Cayley graph over $(\mathbb{Z}/2\mathbb{Z})^{n}$ generated by elements of Hamming weight $n-10\sqrt{n}$. One can check that elements of Hamming weight $\leq n/2-10\sqrt{n}$ form a positive density independent set in this graph. Furthermore, looking at the elements of Hamming weight $n/2-6\sqrt{n}$, this graph has a copy of the Kneser graph $KG(n,n/2-6\sqrt{n})$ and thus has chromatic number at least $\sqrt{n}$, and goes to infinity with $n$. Kříž projects this construction to the integers. See [12] for an exposition of his example by Griesmer and [11] for an extension.
In this paper, we will be concerned with the natural higher-order version of Katznelson’s question. This was first explicitly asked by Huang, Shao, and Ye [15], is the main subject of their book [16], and was asked again as Question 7 in [7]. Before stating it, we need to define topological multiple recurrence.
@@
The set of $n$ so that these generalized polynomials have small $\mathbb{R}/\mathbb{Z}$ is called a “Nil-Bohr” set, or a $Nil_{d}$-Bohr” set when we want to emphasize the parameter of $d$. An ordinary Bohr set is just a $Nil_{1}$-Bohr set.
@@ 2.1 Building the $\ell^{1}$ example
Let $\delta_{1}$ and $\delta_{2}$ be small positive numbers so that $\delta_{1}/\delta_{2}$ is an integer and $\delta_{2}$ is a suitably small function of $\delta_{1}$, so for instance $\delta_{2}<\delta_{1}^{3}$. For example one can pick $\delta_{1}=10^{-2}$ and $\delta_{2}=10^{-10}$. However, for ease of notation we will write $\delta_{1}$ and $\delta_{2}$.
@@
Let $m\in\mathbb{N}\cup\infty$. We say that $B\subset\ell^{1}(\mathbb{N}_{\leq m};\mathbb{R}/\mathbb{Z})$ is a Bohr set if there exists $k$ and $\varepsilon$ and a linear function $f:\ell^{1}(\mathbb{N}_{\leq m};\mathbb{R}/\mathbb{Z})\to(\mathbb{R}/\mathbb{Z})^{k}$ so that $B$ is the set of $b$ so that $|f(b)|<\varepsilon$.
In other words, we must check that for all linear functions $f:\ell^{1}(\mathbb{N};\mathbb{R}/\mathbb{Z})\to(\mathbb{R}/\mathbb{Z})^{k}$, there are points of $S$ that are mapped arbitrarily close to the origin. What we actually need is a finitary version of this.
@@
In fact we only used here that the function is additive, since $\delta_{1}/\delta_{2}\in\mathbb{N}$. The astute reader may notice that the linear functions $f:\ell^{1}(\mathbb{N};\mathbb{R}/\mathbb{Z})\to\mathbb{R}/\mathbb{Z}$ comprise the dual group $\ell^{\infty}(\mathbb{N};\mathbb{R}/\mathbb{Z})$, are given by a sequence of integers $(b_{1},b_{2},\cdots)$ with some $C$ so that $|b_{i}|\leq C$ for all $i$, and send $(a_{1},a_{2},\cdots)$ to $\sum_{i\in\mathbb{N}}a_{i}b_{i}$. In particular, a function to $(\mathbb{R}/\mathbb{Z})^{k}$ will consist of $k$ such functions, one for each coordinate.
@@ 2.2 The coloring
Definition 2.5.
Partition $\mathbb{C}/\mathbb{Z}[i]\cong(\mathbb{R}/\mathbb{Z})^{2}$ into $O(1/\delta^{2}_{2})$ sets of diameter at most $\delta_{2}$. Color $(a_{1},\cdots,a_{m})\in\ell^{1}(\mathbb{N}_{\leq m};\mathbb{R}/\mathbb{Z})$ by the set that contains $\sum_{i=1}^{m}(e(a_{i})-1)$ taken modulo $\mathbb{Z}[i]$.
@@ 3 Concluding Remarks and Open Questions
We generally believe that if Katznelson’s question has a positive answer, the proof somehow relies on topology. Indeed Glasscock, Koutsogiannis, and Richter [7] prove the special case of skew products over $(\mathbb{R}/\mathbb{Z})^{2}$ by using the intermediate value theorem, and [18] relies on the topological lower bound for the chromatic number of the Kneser graph. We also note here that the construction in this paper can be expressed as coming from a skew product over $(\mathbb{R}/\mathbb{Z})^{5}$.
We mention one important obstacle to extending something like our construction to settle the full Katznelson’s question. Let $G$ be an Abelian group. Then an element of $G^{m}$ in the generating set with sum of coordinates 0 can be written in the form $(a_{1}-a_{2},a_{2}-a_{3},\cdots,a_{m-1}-a_{m},a_{m}-a_{1})$, so the cyclic shifts $(a_{1},a_{2},a_{3},\cdots,a_{n})$ and $(a_{2},a_{3},\cdots,a_{n},a_{1})$. Hence, we cannot simply take a function of each coordinate and add it up the way we considered (e(x)1)(e(x)-1) here. This kind of “cyclic shift” obstacle destroys many potential counterexamples to Katznelson’s question, and any kind of high-dimensional projection must break this sort of symmetry. Nonetheless, we believe that our construction represents substantial progress towards Katznelson’s Question.
@@
For an Abelian Cayley graph, we can determine in polynomial time whether there exists a coloring by Bohr sets using <C<C colors by simply checking all such colorings, and in [17] it is proved that Katznelson’s Question is equivalent to the statement that the number of colors necessary in such a coloring is bounded a function of the chromatic number. Hence, if Katznelson’s Question has a positive answer, there is a polynomial algorithm for this “Approximate Graph Coloring” question for Abelian Cayley graphs. For general graphs, such problems are known to be NP-hard conditional on a variant of the Unique Games Conjecture, which gives us reason to believe 3.5. See [5] and the references there for more discussion of Approximate Graph Coloring.
@@
For what sets $S$ does an arbitrary finite coloring of $\mathbb{N}$ always contain a monochromatic $\{x,x+d,x+d^{2}\}$ with $d\in S$?